A diver leaves the springboard with an initial angular velocity of 4.1 rad/sec and a constant angular acceleration of -0.25 rad/sec^2. How long will it take for her to complete 1.5 revolutions, assuming her initial angular displacement is 0 radians?
Answer #1

ozw, t + 2H? 1.5*2* = 4.1 rad xt +x€0.25 mg next t -32. 8t +75.4=0 = 32.81 / 132.8)2 - 4 x.75.4 ta 2 5S +11